OpenStudy (anonymous):

Can someone explain to me how x^2+8x+16 equals (x+4)(x+4)? I don't really understand factoring :\

OpenStudy (anonymous):

Well, because (x+4)(x+4) = x^2 + 4x + 4x + 16 = x^2 + 8x + 16. Distribute the terms and multiply.

OpenStudy (lgbasallote):

notice how if you break down the middle term... x^2 + 4x + 4x + 16 then factor out... x(x+ 4) + 4(x+4) then factor out again... (x+4)(x+4) you have that :D

OpenStudy (lgbasallote):

aww @GT was too fast hahaha

OpenStudy (anonymous):

You can also look at the constant 16 and say what "factors" of 16 add up to 8 (the coefficient of x term). Turns out +4 and +4 do. So, the factors are (x+4) and (x+4).

OpenStudy (anonymous):

Thank you to both :) Your explanations really helped.